OpenStudy (anonymous):

24a^2b^3/4ab^5=? It says to Divide. Write the answer with positive exponents only.

Nnesha (nnesha):

\[\huge\rm \frac{ 24a^2b^3 }{ 4ab^5 }\] apply exponent rule! \[\large \rm\frac{ x^m }{ x^n } = x^{m\color{red}{-}n}\] where base are same at the numerator and at the denominator you can move exponent from bottom to top of the fraction

Nnesha (nnesha):

for example a would be \[a^{2-1}\]

OpenStudy (anonymous):

is it 6a/b^2

Nnesha (nnesha):

:-) perfect!